-- AndreaFestanti - 2016-04-13

Summary of the SPD 2015 data taking

For each period of 2015 data taking the SPD issues and the solutions adopted to fix them are reported.

Main problems during 2015

  1. Wrong strobe length on 8 half-staves

  1. Misbehaviour of "misc" pixel chip internal register
    • first SPD timing calibration (June/July) performed using cosmic rays after the change of the L0-L1 delay allowed to identify the most efficient configuration: hits strobe of 300 ns
    • asymmetric distribution of the IRs discovered in Spetember: the actual position of the hits strobe not centred w.r.t. the collision due to a wrong "misc" value (w.r.t. the most efficient configuration)
    • first test with the correct configuration (misc =128) done on 1st October
      • potential problem in the correlation of the hits and Fast-Or in the data stream spotted by DQM plots
      • back to the previous configuration (misc=192)
    • tests performed to understand the behaviour of the data stream changing the relative position and length of the hits and Fast-Or strobes. Most efficient configuration confirmed to be: 300 ns hits strobe centred w.r.t. the 100 ns Fast-Or strobe
    • using this configuration the correlation between the hits and the Fast-Or in the data-stream is lost. Looking at the temporal matching in the same or previous events:
      • the mismatch appears for all the half-staves after a certain number of events
      • the matching in the same event seems to be periodic and recover after 4 steps
    • TEMPORARY SOLUTION: MEB switched-off, SPD running in SEB mode from 26th to 30th October
    • further set of timing measurements allowed to understand that the problem of the misalignment between hits and Fast-Or information is introduced changing the value of the "misc" parameter when using a 300 ns hits strobe
      • the "misc" is a Pixel Chip internal register that allow to delay the hit data strobe by 100 ns:
        • misc = 128 --> no delay applied
        • misc = 192 --> 100 ns delay applied
      • L0-L1 delay = 6.5 mus during Run1 --> misc = 192
      • L0-L1 delay moved to 7 mus during Run2
    • SOLUTION applied on 30th October starting from run 240854

LHC15a

LHC15b

LHC15c

LHC15d

LHC15e

LHC15f

Run range: 225000-226606

Problem: strobe length larger than standard for 8 HS

Solution: 8 HS masked offline in the OCDB (to be done).

  • The pass2 of LHC15f (August 2015) may still contain problematic chips or hs which had this issue because the OCDB objects were provided before the complete understanding of the problem. The new objects will be provided after finding the solution of this other JIRA ticket https://alice.its.cern.ch/jira/browse/SPD-41*

LHC15g

Run range: 228930-231321

Problem: strobe length larger than standard for 8 HS

Solution: 8 HS masked offline in the OCDB

SPD configuration

Inactive half-staves (masked in the OCDB)

227836-228404 : 4A2, 5A0, 7A0, 8A0, 9A0, 9A1, 9A3, 9A5, 2C0, 2C1, 3C3, 6C0, 7C1, 9C0, 9C3
228930-229245 : 1A0, 1A5, 2A4, 3A2, 4A2, 5A0, 5A3, 7A0, 8A0, 9A0, 9A1, 9A3, 9A5, 2C0, 2C1, 3C3, 6C0, 7C0, 7C1, 9C0, 9C3
229297-230533 : 4A2, 5A0, 7A0, 8A0, 9A0, 9A1, 9A3, 9A5, 2C0, 2C1, 3C3, 6C0, 7C1, 9C0, 9C3
230634-231321 : 4A2, 5A0, 7A0, 8A0, 9A0, 9A1, 9A3, 9A5, 2C0, 2C1, 3C1, 3C3, 6C0, 7C1, 9C0, 9C3

OCDB upload on March 3rd 2016 (https://alice.its.cern.ch/jira/browse/ALIROOT-6569)

LHC15h

Run range: 232914-234050

Problem: strobe length larger than standard for 8 HS

Solution: 8 HS masked offline in the OCDB

SPD configuration

Inactive half-staves (masked in the OCDB)

232914-234062: 4A2, 5A0, 7A0, 8A0, 9A0, 9A1, 9A2, 9A3, 9A5, 2C0, 2C1, 3C1, 3C3, 6C0, 7C1, 9C0, 9C3
235074-235083: 1A5, 4A2, 5A0, 7A0, 8A0, 9A0, 9A1, 9A3, 9A5, 0C5, 2C0, 2C1, 3C1, 3C3, 6C0, 7C1, 9C0, 9C3

OCDB upload on March 18th https://alice.its.cern.ch/jira/browse/ALIROOT-6698

LHC15i

Run range: 235196-235573

Problem: strobe length larger than standard for 8 HS

Solution: 8 HS masked offline in the OCDB

Run range: 235683-236866

Problem: NONE

SPD configuration

Inactive half-staves (masked in the OCDB)

235196-235629: 1A5, 4A2, 5A0, 7A0, 8A0, 9A0, 9A1, 9A3, 9A5, 0C5, 2C0, 2C1, 3C1, 3C3, 6C0, 7C1, 9C0, 9C3, 3A0chip7
235683-236227: 4A2, 5A0, 8A0, 0C5, 2C0, 2C1, 3C1, 3C3, 6C0, 3A0chip7
236234-236248: 4A2, 5A0, 8A0, 0C5, 2C0, 2C1, 3C1, 3C3, 6C0, 3A0chip7, 0A1chip0
236281-236639: 4A2, 5A0, 8A0, 0C5, 2C0, 2C1, 3C1, 3C3, 6C0, 3A0chip7
236862-236866: 1A0, 1A1, 1A2, 1A3, 1A4, 1A5, 4A2, 5A0, 8A0, 0C5, 2C0, 2C1, 3C1, 3C3, 6C0

OCDB upload on March 15th (https://alice.its.cern.ch/jira/browse/ALIROOT-6584)

LHC15j

Run range: 236892-237409

Problem: NONE

Run range: 237502-237505

Problem: Fast-Or and Hits misalignment --> DATA NOT RECOVERABLE

Note: test with misc=128

Run range: 237507-238394

Problem: NONE

Run range: 238395-238433

Problem: Fast-Or and Hits misalignment --> DATA NOT RECOVERABLE

Note: timing scan

Run range: 238451-238621

Problem: NONE

SPD configuration

237115: 9C3

OCDB upload on March 25th (https://alice.its.cern.ch/jira/browse/ALIROOT-6592)

LHC15k

Run range: 238890-239144

Problem: Fast-Or and Hits misalignment --> DATA NOT RECOVERABLE

Note: good runs 238694, 238890

LHC15l

Run range: 239319-240151

Problem: Fast-Or and Hits misalignment -->DATA NOT RECOVERABLE

Note: good runs 239319, 239324, 239270, 239518, 239519, 239655, 239969, 240002, 240069

Run range: 240160-240394

Problem: NONE

Note: SEB mode

Run range: 240398-240440

Problem: Fast-Or and Hits misalignment -->DATA NOT RECOVERABLE

Note: timing scan, good runs 240404, 240427

Run range: 240443-240845

Problem: NONE

Note: SEB mode

Run range: 240854-241544

Problem: NONE

SPD configuration

241187: 0A1

OCDB upload on March 25th (https://alice.its.cern.ch/jira/browse/ALIROOT-6589)

LHC15m

Run range: 243374-244284

Problem: NONE

LHC15n

Run range: 244340-244628

Problem: NONE

LHC15o

Run range: 244824-246994

Problem: NONE

SPD configuration

245731: errors in router 8 --> masked HS 8A3 (OCDB upload on May 18th 2016 https://alice.its.cern.ch/jira/browse/ALIROOT-6699)

246765: errors in router 8 --> only 8 events affected by LinkRx connection loss => not need to mask the hs (8 events out of 800k).

Edit | Attach | Watch | Print version | History: r8 < r7 < r6 < r5 < r4 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r8 - 2016-10-11 - AnnalisaMastroserio
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    AliceSPD All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback